OpenStudy (anonymous):

3|4w - 1| - 5 = 10 I got w = 6/4 or w = -1 and when I input them back both work, the back of the book says it's just -1. Am I doing something wrong or is the book wrong?

jimthompson5910 (jim_thompson5910):

you should get two solutions and they are w = 3/2 or w = -1

jimthompson5910 (jim_thompson5910):

so the fact that 3/2 is left out is a typo

OpenStudy (anonymous):

Yeah I have that there, 6/4 = 3/2.

jimthompson5910 (jim_thompson5910):

so it looks like the book is wrong I guess if they wanted integer only solutions, then that would mean that you exclude 3/2

OpenStudy (anonymous):

Ok thanks a lot!

jimthompson5910 (jim_thompson5910):

yw